# Sylva: Tailoring Personalized Adversarial Defense in Pre-trained Models via Collaborative Fine-tuning

### v1

@@
Abstract.
The growing adoption of large pre-trained models in edge computing has made deploying model inference on mobile clients both practical and popular. These devices are inherently vulnerable to direct adversarial attacks, which pose a substantial threat to the robustness and security of deployed models. Federated adversarial training (FAT) has emerged as an effective solution to enhance model robustness while preserving client privacy. However, FAT frequently produces a generalized global model, which struggles to address the diverse and heterogeneous data distributions across clients, resulting in insufficiently personalized performance, while also encountering substantial communication challenges during the training process.
In this paper, we propose Sylva, a personalized collaborative adversarial training framework designed to deliver customized defense models for each client through a two-phase process. In Phase 1, Sylva employs LoRA for local adversarial fine-tuning, enabling clients to personalize model robustness while drastically reducing communication costs by uploading only LoRA parameters during federated aggregation. In Phase 2, a game-based layer selection strategy is introduced to enhance accuracy on benign data, further refining the personalized model. This approach ensures that each client receives a tailored defense model that balances robustness and accuracy effectively. Extensive experiments on benchmark datasets demonstrate that Sylva can achieve up to 50$\times$ improvements in communication efficiency compared to state-of-the-art algorithms, while achieving up to 29.5% and 50.4% enhancements in adversarial robustness and benign accuracy, respectively.
Pre-trained Models, Personalized federated learning, Adversarial training, Fine-tuning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13-17, 2025; Taipei, Taiwanisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacyccs: Computing methodologies Artificial intelligence
1. Introduction
@@
With the rapid advancement of large language models (LLM), large-scale pre-trained models have garnered widespread attention across various fields, including computer vision (Dosovitskiy et al., 2021), autonomous driving (Yang et al., 2024), and healthcare (Jumper et al., 2021), etc. Fine-tuning pre-trained models for downstream tasks has gradually established itself as a novel learning paradigm (Houlsby et al., 2019; Hu et al., 2022; Liu et al., 2022a). Meanwhile, the increasing computational power of edge devices has facilitated the localized deployment of the pre-trained models, unlocking their potential for various applications on devices (Liu et al., 2024; Yuan et al., 2024).
@@
However, recent studies have revealed substantial security risks associated with deploying pre-trained models on edge devices. The typically weak protection mechanisms of these devices render their various permissions susceptible to unauthorized access, exposing sensitive resources to potential exploitation (Zhou et al., 2021). Furthermore, the parameters of the machine learning model stored locally on edge devices are often inadequately secured, increasing the risk of theft or manipulation (Deng et al., 2022). On the other hand, many pre-trained model weights and architectures are open-source, allowing attackers to easily obtain original parameters and designs from public repositories. This accessibility enables the development of more sophisticated attack strategies, such as embedding backdoors or executing poisoning attacks on fine-tuned models for downstream tasks (Jia et al., 2022; Liu et al., 2022b). Such vulnerabilities pose significant threats to the reliability and security of edge-based machine learning applications, highlighting an urgent need for robust defense mechanisms.
@@
To mitigate these challenges, adversarial training (AT) has become a widely utilized defense strategy (Madry et al., 2018; Zhang et al., 2019; Lee et al., 2020). By incorporating benign and adversarial data during the training process, this approach enhances model robustness while striving to maintain accuracy. Despite its effectiveness, traditional adversarial training faces notable limitations when applied in real-world scenarios. As shown on the left side of Fig. 1, the challenges of adversarial training stem from several key factors. First, the limited data available to local clients often hinders effective adversarial training, leading to suboptimal performance or increased risks of overfitting during fine-tuning for downstream tasks. Second, the data across different clients is non-independent and identically distributed (non-IID). For instance, in autonomous driving, data collected under extreme weather conditions is scarce (Marathe et al., 2023), while in healthcare, datasets from different hospitals exhibit significant variability (Xie et al., 2024). This variability exposes weaknesses that attackers can exploit, such as rare weather conditions in autonomous driving or underrepresented patient groups in healthcare. Additionally, privacy concerns prohibit uploading client data to the cloud for centralized training, further exacerbating these challenges.
@@
Figure 1. Challenges of adversarial training in multi-client scenario. (1) Left: Local traditional adversarial training performs well for classes with abundant data but struggles with those that are rare or unseen. (2) Right: Federated adversarial training addresses these issues by improving generalization across clients. However, it lacks the necessary personalization to meet the specific needs of individual clients and their unique data distributions.
@@
In multi-client environments, federated adversarial training (FAT) has proven to be a powerful approach for enhancing model robustness in multi-client environments while preserving privacy (Zizzo et al., 2020). As large pre-trained models gain prominence, fine-tuning through FAT is expected to become a dominant trend, enabling devices to contribute their limited data via periodic synchronization and knowledge sharing. Despite its potential, most existing studies have concentrated on building generalized global models to address the challenges of non-IID data (Zhou et al., 2022; Chen et al., 2022; Li et al., 2023a). However, in the presence of significant data heterogeneity, such global models often fail to deliver effective defense tailored to individual clients. Their strong generalization, while advantageous in some contexts, limits their ability to adapt to unique client-specific data distributions, resulting in diminished performance, as illustrated on the right side of Fig. 1. Additionally, the integration of FAT with large pre-trained models introduces communication overhead, which may significantly impact training speed, particularly in resource-constrained environments. These limitations expose a glaring gap in current methodologies and raise an urgent question: Can we design a personalized adversarial training framework that provides robust and tailored defense models for each client while preserving the training efficiency on edge devices?
@@
This study aims to address three critical challenges in adversarial fine-tuning pre-trained models in multi-client environments: (1) Designing a personalized defense framework: Given the heterogeneous data distributions across clients, a collaborative adversarial training framework is required to provide each client with a tailored defense model while ensuring data privacy. (2) Enhancing both adversarial robustness and benign accuracy: Developing a training paradigm to improve the adversarial robustness of models while maintaining high accuracy on benign data is crucial to ensure reliable performance in diverse and challenging scenarios. (3) Balancing efficiency for clients: Achieving improved performance on resource-constrained edge devices requires addressing communication and training efficiency challenges. By achieving these objectives, each client can obtain a model that is both robust and highly accurate, tailored specifically to its unique data distribution. Such models excel in downstream tasks under typical conditions while providing strong, personalized defenses against adversarial attacks targeting individual clients.
@@
To address the challenges above, we propose Sylva, a collaborative framework that tailors pre-trained models with robust defense capabilities through personalized adversarial fine-tuning for edge devices. Sylva is designed to fine-tune models in a way that simultaneously enhances robustness and accuracy for each client, even in environments with heterogeneous data distributions. To validate the feasibility of our approach, we first conduct preliminary experiments leveraging the low-rank adaptation (LoRA) algorithm for personalized adversarial fine-tuning in real-world scenarios. These experiments not only demonstrate the potential of LoRA in adversarial training but also expose the critical limitations of existing defense strategies. Then we present Sylva’s two-phase training framework. In the first phase, personalized adversarial training divides the model into two modules: the LoRA module and the classifier module. The parameters of the LoRA module are shared and aggregated across clients to enhance the generalization ability of the model backbone for adversarial feature extraction. Meanwhile, the classifier module is fine-tuned locally on each client, enabling personalized improvements in classifying adversarial samples based on each client’s unique data distribution. We propose a novel adaptive class-balanced dynamic weighted loss for handling data heterogeneity and a ball-tree-based aggregation algorithm to accelerate adversarial training convergence. In the second phase, we employ a Shapley value-based cooperative game approach to enhance the accuracy of benign data for each client. We identify and selectively freeze the optimal layers within each model, striking a critical balance that maximizes accuracy of benign data while minimizing any compromise in robustness. This strategic layer adjustment ensures that each client’s model achieves a higher degree of personalization, effectively navigating the trade-off between benign accuracy and adversarial robustness. The main contributions of this paper can be summarized as follows:
@@
  • To our best knowledge, we propose the first personalized collaborative adversarial training framework for pre-trained models, enabling client-specific defenses while preserving privacy and minimizing overhead on edge devices.
@@
  • We introduce a novel training paradigm for LoRA-based adversarial fine-tuning, integrating adaptive optimization strategies with a ball-tree-based model aggregation and a newly designed loss, to enhance robust generalization under heterogeneous data distributions.
  • We propose a cooperative game-based layer freezing method for pre-trained models, leveraging a novel value function and Monte Carlo sampling to optimize the trade-off between benign data accuracy and model robustness.
@@
  • We conduct experiments on widely used datasets and pre-trained models, comparing Sylva with popular defenses under various attack scenarios, demonstrating its effectiveness in personalized defense.
2. Preliminaries
2.1. Federated Adversarial Training
@@
Federated learning is an effective training algorithm for protecting client data privacy (McMahan et al., 2017; Qi et al., 2025; Liu et al., 2022c). In a distributed system with $N$ clients and a central server, each client $i$ has its private dataset $(x,y)\in\mathcal{D}_{i}$, where $x$ represents the benign samples, and $y$ denotes the ground truth. The server initializes the model, and client $i$ downloads the model, denoted as $w_{i}$ and trains it using local data. The training objective for each client can be expressed as
(1) $f_{i}(w_{i})=\frac{1}{|\mathcal{D}_{i}|}\sum_{(x,y)\in\mathcal{D}_{i}}f_{i}(w_{i},x,y),$
@@
For federated adversarial training, the local training objective is to enable the model to recognize both benign samples and adversarial samples, thereby improving its robustness through adversarial training. For a benign sample $x$, we add noise $\sigma$ to generate an adversarial sample $x^{adv}=x+\sigma$. The goal of generating robust adversarial samples is to find the noise that maximizes the model loss $\mathcal{L}(w_{i},x+\sigma,y)$. The local training objective for clients in federated adversarial training can be expressed as
(2) $f_{i}(w_{i},x,x^{adv},y)=\min\mathbb{E}_{x\sim\mathcal{D}_{i}}\left[\max_{\left\|x^{adv}-x\right\|_{\infty}\leq\delta}\mathcal{L}(w_{i},x,x^{adv},y)\right].$
After training locally for a certain number of rounds, clients upload their models to the server, where parameter aggregation is performed using the data size of each device as a weight, represented as
(3) $w_{g}=\sum_{i=1}^{N}\frac{|\mathcal{D}_{i}|w_{i}}{\sum_{i=1}^{N}|\mathcal{D}_{i}|}.$
Subsequently, clients can download the global model $w_{g}$ for further local training. This process is repeated until final convergence.
2.2. Low-Rank Adaptation
Fine-tuning pre-trained models in mobile computing is challenging due to limited GPU memory, which often makes training all model parameters impractical. Low-rank adaptation (LoRA) (Hu et al., 2022) addresses this issue by freezing pre-trained model parameters and introducing lightweight trainable layers into transformer modules, significantly reducing computational overhead.
During the fine-tuning process, the updated parameters exhibit a low intrinsic rank. For a LoRA fine-tuned model $w\in\mathbb{R}^{r_{in}\times C}$, where $r_{in}$ represents the input dimension, and $C$ denotes the total number of prediction classes. The entire model can be divided into two parts: the backbone $w^{B}\in\mathbb{R}^{r_{in}\times r_{out}}$ and the classifier $w^{C}\in\mathbb{R}^{r_{out}\times C}$. The backbone’s parameters can be represented as the combination of the pre-trained model parameters $w^{P}\in\mathbb{R}^{r_{in}\times r_{out}}$ and the LoRA module parameters $w^{L}\in\mathbb{R}^{r_{in}\times r_{out}}$. The LoRA parameters $w^{L}$ can be expressed as a product of two low-rank matrices $w^{LA}\in\mathbb{R}^{r_{in}\times r}$ and $w^{LB}\in\mathbb{R}^{r\times r_{out}}$. We define the overall mapping of the model as $\mathcal{F}(\cdot)$, where the pre-trained model mapping and LoRA module mapping are $\mathcal{F}^{P}(\cdot)$ and $\mathcal{F}^{L}(\cdot)$, and the classifier mapping is $\mathcal{F}^{C}(\cdot)$. For a sample $x$ the model output can be expressed as
(4) $\hat{y}=\mathcal{F}(x)=\mathcal{F}^{C}(\mathcal{F}^{P}(x)+\mathcal{F}^{L}(x)).$
During training, the parameters of $w^{P}$ remain frozen, and only $w^{L}$ is updated, achieving a performance comparable to full-parameter fine-tuning.
2.3. Threat Model
@@
2.3.1. Attacker
We posit that deploying models on client devices introduces potential system-level vulnerabilities, which attackers can exploit to gain access to client permissions and compromise the model’s inference process through adversarial attacks.
Depending on the level of access an attacker has to the client model, these attacks can be classified into two categories:
  • White-box attacks. The attacker has full access to all parameters of the deployed model and can leverage this information to generate highly effective adversarial samples. This scenario forms the basis for most existing attack methodologies.
  • Grey-box attacks. The attacker knows the pre-trained model from which the client’s task is fine-tuned and can access the original pre-trained model parameters (e.g., publicly platforms like Hugging Face) but lacks access to the fine-tuned parameters deployed on the client.
@@ 2.3.1. Attacker
The objective of the attacker in both scenarios is to exploit the accessible model parameters to execute adversarial attacks on the client’s downstream tasks. The adversarial samples generated must meet two critical criteria: (1) Inconspicuousness. The perturbations in the adversarial samples should be subtle enough to go unnoticed by humans. (2) Impactfulness. The adversarial attacks must effectively compromise the model’s downstream tasks, even when standard defense mechanisms are in place.
2.3.2. Defender
@@
We propose involving defenders in the fine-tuning process of client models for downstream tasks to enhance robustness. Collaborative fine-tuning is assumed to take place within a distributed framework based on a parameter server (PS) architecture (Li et al., 2014). In this setup, the defender operates at the server level, providing macro-level oversight and regulation of client activities to ensure robust and secure model training.
Defenders should possess the following three characteristics:
  • Data Privacy. The defense mechanism must safeguard the privacy and security of each client’s private data, ensuring no compromise occurs during the training or deployment process.
  • Heterogeneity. Clients operate in diverse scenarios, requiring the defense approach to support personalized models that account for individual preferences and adversarial capabilities.
  • Efficiency. To accommodate the limited computational resources of clients and mitigate potential network congestion, the defense strategy should prioritize efficiency, minimizing both memory usage and communication overhead.
In the FAT scenario, it is considered that for clients with similar data distributions, there are dedicated and advanced secure devices specifically used for data collection and defense training. For example, in autonomous driving, secure vehicles can be utilized to collect driving data and train defense models, ensuring robust security during the training process. Upon completing the training phase, the models are deployed to other vehicles with similar distribution but lower security capabilities, such as those within the same geographic region, to perform inference tasks. The FAT training paradigm ensures that attackers can only target the inference process on deployed devices, as the secure vehicles used during the training phase, equipped with robust defenses, remain inaccessible and protected from attacks.
Table 1. Adversarial training with/without LoRA
@@
Figure 2. Performance of personalized LoRA in adversarial training under heterogeneous environments
@@
Figure 3. PCA embedding of the LoRA model after federated adversarial training
2.4. Preliminary Analysis of Existing Issues
In this section, we explore the critical issues in fine-tuning pre-trained models for multi-client attack-defense systems. To address these issues, we design experiments to validate their impact, offering essential insights for the development of Sylva. Specifically, we raise four key research questions:
  • Q1: How can the computational resources of edge devices be effectively leveraged for adversarial defense training?
  • Q2: How can personalized framework be tailored to meet the unique requirements of each client?
  • Q3: How can valuable knowledge be effectively extracted from clients with diverse data distributions?
  • Q4: How can the trade-off between adversarial robustness and benign accuracy be systematically optimized?
We conduct preliminary experiments to explore the above questions, highlighting our discoveries for each experiment and identifying remaining challenges that guide the design of Sylva.
2.4.1. Impact of Adversarial Training with LoRA
Adversarial training via full model training has demonstrated effectiveness, but it incurs high training costs, rendering it inefficient for edge devices with limited computational resources. To address this, we conduct experiments to evaluate the effectiveness of adversarial fine-tuning using LoRA.
We utilize the widely-used ResNet (He et al., 2016) and ViT (Dosovitskiy et al., 2021) (ViT-T/16, ViT-B/16, ViT-L/16) pre-trained models, applying the TRADES algorithm (Zhang et al., 2019) for adversarial fine-tuning via LoRA on the CIFAR-10 dataset (Krizhevsky et al., 2009). ResNet and ViT-T are trained for 50 epochs, while ViT-B and ViT-L are trained for 100 epochs. The test environment is simulated according to the setup described in Section 4. The experimental results are evaluated using five metrics: Adversarial Robustness (AR), assessed by accuracy on adversarial samples; Benign Accuracy (BA), assessed by accuracy on benign samples; Time, which tracks the time required per training epoch; Parameter Size (Paras), which quantifies the number of model parameters fine-tuned during training; GPU Memory (Mem), the GPU required for adversarial training with a fixed batch size.
The experimental results, presented in Table 1, demonstrate that adversarial fine-tuning using LoRA significantly reduces training time by up to 23.2% compared to full model fine-tuning. Moreover, LoRA requires only 0.7% of the parameters needed for full model training, saving approximately 30.5% of GPU memory during training and making it highly efficient in resource-constrained edge devices. Additionally, as model size increases, both AR and BA improve with LoRA-based adversarial fine-tuning. This aligns with findings in other tasks (Hu et al., 2022; Savostianova et al., 2024), where as the model size increases, the low-rank eigenvalues in the model’s mapping space capture most of the energy. Consequently, low-rank fine-tuning enhances adaptation to downstream tasks while preserving the model’s overall feature extraction capabilities.
Discovery 1: LoRA-based adversarial fine-tuning preserves or even improves defense effectiveness while drastically reducing computational overhead, making it ideal for edge devices.
However, a critical challenge arises when scaling from single-client to multi-client training, as it requires designing a collaborative adversarial defense framework using LoRA to fully leverage its advantages.
2.4.2. Impact of the Personalized Framework
Given the powerful feature extraction capabilities of pre-trained model backbones, we believe that their feature extraction ability remains effective for adversarial examples after adversarial fine-tuning. To explore this, we conduct an experiment simulating two clients, where the training and testing data for each client are split according to the same Dirichlet distribution, thereby simulating a unique non-IID environment for each client (Li et al., 2020a). We utilize three different ViT models and implement a federated adversarial training strategy, applying TRADES for local fine-tuning through LoRA, on the CIFAR-10 (Krizhevsky et al., 2009), STL-10 (Coates et al., 2011), and GTSRB (Stallkamp et al., 2011) datasets. The federated adversarial training involve 5 local training epochs followed by parameter upload and cloud aggregation, repeated for 15 rounds. For the model upload step, we compare two strategies: uploading whole parameters (W) versus uploading only the LoRA parameters of the backbone while excluding the classifier parameters (L). The latter approach enables the training of a shared global backbone while allowing each client to personalize its classifier.
Discovery 2: Aggregating only the backbone (LoRA) and personalizing the classifier preserves generalized feature extraction while enhancing the classifier’s sensitivity to heterogeneity, ensuring both global generalization and personalized defense.
@@ 2.4.2. Impact of the Personalized Framework
Figure 4. Comparison of aggregation methods for adversarial training in heterogeneous environments

### v3

# Lorica: A Synergistic Fine-Tuning Framework for Advancing Personalized Adversarial Robustness

@@
Abstract
[…]
I Introduction
@@
With the rapid advancement of large language models (LLM), large-scale pre-trained models have garnered widespread attention across various fields, including computer vision [9] and autonomous driving [39], etc. Fine-tuning pre-trained models for downstream tasks has gradually established itself as a novel learning paradigm [14]. Meanwhile, the increasing computational power of edge devices has facilitated the localized deployment of the pre-trained models, unlocking their potential for various applications on devices [22].
@@
However, recent studies have revealed substantial security risks associated with deploying pre-trained models on edge devices. The weak protection mechanisms of these devices make permissions vulnerable to unauthorized access, exposing sensitive resources, while locally stored model parameters are often inadequately secured, heightening the risk of theft or manipulation [7]. On the other hand, many pre-trained model weights and architectures are open-source, allowing attackers to easily obtain original parameters and designs from public repositories, such as embedding backdoors or executing poisoning attacks on fine-tuned models for downstream tasks [20]. These vulnerabilities threaten the security of edge-based learning, highlighting the need for robust defenses.
@@
To mitigate these challenges, adversarial training (AT) has become a widely utilized defense strategy [25, 42]. By incorporating benign and adversarial data, this approach enhances model robustness while preserving accuracy, yet traditional adversarial training shows notable limitations in real-world scenarios. As shown on the left side of fig. 1, the challenges of adversarial training stem from several key factors. First, the limited data of local clients often hinders adversarial training, causing suboptimal performance or overfitting during downstream fine-tuning. Second, client data is typically non-IID; for example, autonomous driving data under extreme weather is scarce, while healthcare datasets vary significantly across hospitals. Such variability exposes weaknesses exploitable by attackers, and privacy concerns further prevent centralized training, exacerbating these challenges.
@@
Figure 1: Challenges of adversarial training in multi-client scenario. (1) Left: Local adversarial training performs well on abundant classes but fails on rare or unseen ones. (2) Right: Federated adversarial training improves generalization across clients but lacks personalization for unique client data distributions.
@@
In multi-client environments, federated adversarial training (FAT) has proven effective for enhancing robustness while preserving privacy [49]. With large pre-trained models gaining prominence, FAT-based fine-tuning is expected to dominate, allowing devices to contribute limited data through periodic synchronization and knowledge sharing. Despite its potential, most studies focus on generalized global models for non-IID data [46, 4]. Yet under significant heterogeneity, such models often fail to provide effective client-specific defense, as their strong generalization limits adaptation to unique data distributions, leading to diminished performance, as illustrated on the right side of fig. 1. Moreover, FAT typically assumes benign participants. Under Byzantine clients uploading arbitrary updates, plug-in defenses can help, but integrating them into the FAT pipeline often adds extra computation overhead in large model training [24]. Integrating FAT with large pre-trained models introduces communication overhead that can hinder training speed, especially in resource-constrained settings. These limitations highlight a critical gap: can we design a personalized adversarial training framework that delivers robust, tailored defense for each client while maintaining efficiency on edge devices and simultaneously being resilient to Byzantine attacks?
@@
This study aims to address three critical challenges in adversarial fine-tuning pre-trained models in multi-client environments: (1) Designing a personalized defense framework: Given the heterogeneous data distributions across clients, a collaborative adversarial training framework is required to provide each client with a tailored defense model while ensuring data privacy. (2) Enhancing robustness under adversarial and Byzantine attacks: Ensuring high benign accuracy and adversarial robustness in the presence of Byzantine clients uploading malicious updates is critical for reliable deployment. (3) Balancing efficiency for clients: Achieving improved performance on resource-constrained edge devices requires addressing communication and training efficiency challenges. By achieving these objectives, each client obtains a robust and accurate model tailored to its unique data, excelling in downstream tasks while offering strong personalized defenses against adversarial attacks.
@@
To address the challenges above, we propose Lorica, a synergistic framework that tailors pre-trained models with robust defense capabilities through personalized adversarial fine-tuning for edge devices. Lorica is designed to fine-tune models in a way that simultaneously enhances robustness and accuracy for each client, even in environments with heterogeneous data distributions. To validate feasibility, we conduct preliminary experiments using the memory-efficient low-rank adaptation (LoRA-FA) algorithm for personalized adversarial fine-tuning. These experiments demonstrate its potential in adversarial training while exposing critical limitations of existing defenses. Then we present Lorica’s two-phase training framework. In the first phase, personalized adversarial training divides the model into two modules: the LoRA-FA module and the classifier module. The parameters of the LoRA-FA module are shared and aggregated across clients to enhance the generalization ability of the model backbone for adversarial feature extraction. Meanwhile, the classifier module is fine-tuned locally on each client, enabling personalized improvements in classifying adversarial samples based on each client’s unique data distribution. We propose a ball-tree-based aggregation algorithm to accelerate adversarial training convergence and a MoE-inspired adaptive class-balanced weighted loss to address data heterogeneity. Meanwhile, the aggregation also provides Byzantine awareness by mitigating the influence of malicious client updates. In the second phase, we employ a forward-gating network to adaptively select and freeze layers, enhancing each client’s benign accuracy while preserving robustness. This selective adjustment enables a higher degree of personalization, effectively balancing benign performance with adversarial defense. The main contributions of this paper can be summarized as follows:
@@
  • We propose a personalized synergistic adversarial training framework for pre-trained models, enabling client-specific defenses while preserving privacy and minimizing overhead on edge devices.
@@
  • We introduce LoRA-FA adversarial fine-tuning, integrating MoE-inspired adaptive optimization with a ball-tree-based model aggregation with Byzantine-aware and a newly designed loss to enhance robust generalization under heterogeneous data, along with theoretical convergence guarantees.
  • We propose a cooperative a forward-gating layer selection method for pre-trained models, where the gating network adaptively selects layers to optimize the trade-off between benign accuracy and adversarial robustness.
@@
  • We conduct experiments on widely used datasets and pre-trained models, comparing Lorica with popular defenses under various attack scenarios, demonstrating its effectiveness in personalized defense.
II Preliminaries
II-A Federated Adversarial Training
@@
Federated learning is an effective algorithm for protecting client privacy [26, 31]. In a distributed system with $N$ clients and a server, each client $i$ holds a private dataset $(x,y)\in\mathcal{D}_{i}$, downloads the initialized model $w_{i}$, and trains it on local data. The training objective for each client can be expressed as
$f_{i}(w_{i})=\frac{1}{|\mathcal{D}_{i}|}\sum_{(x,y)\in\mathcal{D}_{i}}f_{i}(w_{i},x,y),$ (1)
@@
In federated adversarial training, the local objective is to train models to recognize both benign and adversarial samples, thereby improving robustness. For a benign sample $x$, an adversarial sample is generated as $x^{adv}=x+\sigma$, where the noise $\sigma$ maximizes the loss $\mathcal{L}(w_{i},x+\sigma,y)$. The local training objective for clients in federated adversarial training can be expressed as
$f_{i}(w_{i},x,x^{adv},y)=\min\mathbb{E}_{x\sim\mathcal{D}_{i}}\left[\max\mathcal{L}(w_{i},x,x^{adv},y)\right].$ (2)
After local training, clients upload their models to the server, which aggregates parameters weighted by the data size of each device, represented as
$w_{g}=\sum_{i=1}^{N}\frac{|\mathcal{D}_{i}|w_{i}}{\sum_{i=1}^{N}|\mathcal{D}_{i}|}.$ (3)
Subsequently, clients can download the global model $w_{g}$ for further local training, repeating until final convergence.
Most FAT methods assume benign clients during adversarial training, yet the training process itself is vulnerable to Byzantine behaviors. We consider a threat model where a fraction $\rho$ of clients are malicious and can upload arbitrary LoRA-FA updates (or parameters) to disrupt global or expert aggregation, degrading robustness and benign accuracy. The set of honest clients is denoted as $\mathbf{H}\subseteq\{1,2,\dots,N\}$.
II-B Memory-efficient Low-Rank Adaptation
Fine-tuning pre-trained models in mobile computing is challenging due to limited GPU memory, which makes full-parameter training impractical. LoRA-FA [44], built on low-rank adaptation (LoRA) [14], addresses this issue by freezing pre-trained parameters and introducing lightweight trainable layers into transformer modules, thereby reducing computational overhead.
During fine-tuning, the updated parameters exhibit low intrinsic rank. For a LoRA-FA model $w\in\mathbb{R}^{r_{in}\times C}$, where $r_{in}$ is the input dimension and $C$ the number of classes, the model consists of a backbone $w^{B}\in\mathbb{R}^{r_{in}\times r_{out}}$ and a classifier $w^{C}\in\mathbb{R}^{r_{out}\times C}$. The backbone is composed of frozen pre-trained parameters $w^{P}$ and LoRA-FA parameters $w^{L}$, with $w^{L}=w^{LA}w^{LB}$, where $w^{LA}$ is fixed and only $w^{LB}$ is updated. We define the overall mapping as $\mathcal{F}(\cdot)$, where $\mathcal{F}^{P}(\cdot)$, $\mathcal{F}^{L}(\cdot)$, and $\mathcal{F}^{C}(\cdot)$ denote the pre-trained backbone, LoRA-FA module, and classifier, respectively. For an input $x$, the model output is
$\hat{y}=\mathcal{F}(x)=\mathcal{F}^{C}(\mathcal{F}^{P}(x)+\mathcal{F}^{L}(x)).$ (4)
During training, $w^{P}$ and $w^{LA}$ remain frozen, while only $w^{LB}$ is updated, achieving efficiency with performance close to full-parameter fine-tuning.
II-C Threat Model
@@
II-C1 Attacker
@@
We posit that deploying models on client devices introduces system-level vulnerabilities, allowing attackers to exploit client permissions and compromise model inference through adversarial attacks. Depending on the attacker’s access level, such attacks can be classified into two categories: (1) White-box attacks. The attacker has full access to the deployed model parameters and can generate strong adversarial examples. (2) Grey-box attacks. The attacker has access to the pre-trained model (e.g., from public platforms such as Hugging Face) but not the client-specific fine-tuned parameters. The adversarial samples generated must meet two critical criteria: (1) Inconspicuousness. The perturbations in the adversarial samples should be subtle enough to go unnoticed by humans. (2) Impactfulness. The adversarial attacks must effectively compromise the model’s downstream tasks, even when standard defense mechanisms are in place.
In addition to inference-time adversaries, we also consider Byzantine attackers during federated adversarial training, a setting rarely explored in existing FAT studies. We assume that a fraction $\rho$ of clients can behave maliciously and upload arbitrary LoRA-FA updates or parameters to the server, aiming to disrupt model aggregation and degrade robustness and benign accuracy.
II-C2 Defender
@@
We propose involving defenders in the fine-tuning process of client models for downstream tasks to enhance robustness. Collaborative fine-tuning is assumed to take place within a distributed framework based on a parameter server (PS) architecture [16]. In this setup, the defender operates at the server level, providing macro-level oversight and regulation of client activities to ensure robust and secure model training.
Defenders should satisfy three key properties: (1) Data Privacy, ensuring that each client’s private data remains protected throughout training and deployment; (2) Heterogeneity, supporting personalized defenses that adapt to diverse client data distributions, adversarial capabilities, and the presence of Byzantine or malicious participants; and (3) Efficiency, minimizing computation, memory usage, and communication overhead to enable deployment on resource-constrained devices and networks.
In contrast to conventional FAT settings that focus on inference-time adversarial attacks, we consider a defender operating during federated adversarial training under heterogeneous client data distributions. The primary goal remains effective coordination of adversarial training for personalized robustness, while also accounting for possible Byzantine behaviors arising from a small fraction of clients during training.
TABLE I: Adversarial training with/without LoRA and LoRA-FA
@@
Figure 2: Performance of personalized LoRA-FA in adversarial training under heterogeneous environments
@@
Figure 3: PCA embedding of the LoRA-FA after federated adversarial training
@@
Figure 4: Comparison of aggregation methods for adversarial training in heterogeneous environments
II-D Preliminary Analysis of Existing Issues
In this section, we investigate the critical challenges of fine-tuning pre-trained models for multi-client attack-defense systems, including the efficient use of edge resources for adversarial training, the need for personalized frameworks, the extraction of knowledge from heterogeneous data, and the trade-off between robustness and benign accuracy. We design preliminary experiments to examine these factors, highlighting key findings and open challenges that motivate the development of Lorica.
II-D1 Impact of Adversarial Training with LoRA-FA
Adversarial training via full model training has demonstrated effectiveness, but it incurs high training costs, rendering it inefficient for edge devices with limited computational resources. To address this, we conduct experiments to evaluate the effectiveness of adversarial fine-tuning using LoRA-FA.
We utilize the widely-used ResNet [12] and ViT [9] pre-trained models, applying the TRADES algorithm [42] for adversarial fine-tuning via LoRA-FA on the CIFAR-10 dataset [15]. ResNet and ViT-T are trained for 50 epochs, while ViT-B and ViT-L are trained for 100 epochs. The experimental results are evaluated using five metrics: Adversarial Robustness (AR), assessed by accuracy on adversarial samples; Benign Accuracy (BA), assessed by accuracy on benign samples; Time, which tracks the time required per training epoch; Parameter Size (Paras), which quantifies the number of model parameters fine-tuned during training; GPU Memory (Mem), the GPU required for adversarial training with a fixed batch size.
The results in table I show that adversarial fine-tuning with LoRA-FA reduces training time by up to 31.2% compared to full-model fine-tuning, while requiring only 0.4% of the parameters and saving about 70.8% of GPU memory, making it highly efficient for edge devices. As model size increases, LoRA-based adversarial fine-tuning improves both AR and BA, with LoRA-FA achieving comparable performance at lower computational cost. This is consistent with prior findings [14] that low-rank eigenvalues capture most of the model’s energy and preserve feature extraction.
Discovery 1: LoRA-based adversarial fine-tuning preserves or even improves defense effectiveness while drastically reducing computational overhead.
However, a critical challenge arises when scaling from single-client to multi-client training, as it requires designing a collaborative adversarial defense framework using LoRA-FA to fully leverage its advantages.
II-D2 Impact of the Personalized Framework
Given the powerful feature extraction capabilities of pre-trained model backbones, we believe that their feature extraction ability remains effective for adversarial examples after adversarial fine-tuning. To explore this, we conduct an experiment, where the training and testing data for each client are split according to the same Dirichlet distribution, thereby simulating a unique non-IID environment for each client. We utilize three different ViT models and implement a federated adversarial training strategy, applying TRADES for local fine-tuning through LoRA-FA, on the CIFAR-10 [15], STL-10 [5], and GTSRB [34] datasets. The federated adversarial training involve 5 local training epochs followed by parameter upload and cloud aggregation, repeated for 15 rounds. For the model upload step, we compare two strategies: uploading whole parameters (W) versus uploading only the LoRA-FA parameters of the backbone while excluding the classifier parameters (L). The latter approach enables the training of a shared global backbone while allowing each client to personalize its classifier.
Discovery 2: Aggregating only the backbone (LoRA-FA) and personalizing the classifier preserves generalized feature extraction while enhancing the classifier’s sensitivity to heterogeneity, ensuring both global generalization and personalized defense.